40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11978     Entry time: Fri Feb 5 15:02:13 2016     In reply to: 11968
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: X-end NE cable 

[Steve, gautam]

Steve thinks that the X-end Innolight does not come with the noise-eater option (it is an add-on and not a standard feature, and the purchase order for the PSL Innolight explicitly mentions that it comes with the NE option, but the X-end Innolight has no such remarks), which would explain why there is no difference with the noise eater ON/OFF. During earlier investigations however, I had found that there was a cable labelled "Noise-Eater" connected to one of the Modulation Inputs on the rear of the Innolight controller. Today, we traced this down. The modulation input on the rear says "Current Laser Diode 0.1A/V". To this input, a Tee is connected, one end of which is terminated with a 50ohm terminator. The other end of the Tee is connected to a BNC cable labelled "Nosie-Eater", which we traced all the way to the PSL table, where it is just hanging (also labelled "X end green noise eater"), unterminated, at the southeast corner of the PSL table. It is unlikely that this is of any consequence given the indicated coefficient of 0.1A/V, but could this somehow be introducing some junk into the laser diode current which is then showing up as intensity fluctuations in the output? Unfortunately, during the PLL measurements, I did not think to disconnect this BNC and take a spectrum. It would also seem that the noise-eater feedback to the laser diode current is implemented internally, and not via this external modulation input jack (the PSL, which I believe has the noise-eater enabled, has nothing connected to this rear input)...

 

ELOG V3.1.3-